OpenStudy (anonymous):

question 1 3a - 4 ≤ 5 =? explain how you got the answer question 2 n/-3 + 5 > 4 explain how you got the answer @zepdrix @pooja

OpenStudy (anonymous):

@Kitten_is_back @imqwerty @Agl202

OpenStudy (alivejeremy):

do u go to k12

OpenStudy (anonymous):

what do you mean

OpenStudy (alivejeremy):

nvm

OpenStudy (anonymous):

i get home schooled if that what u mean

pooja195 (pooja195):

3a - 4 ≤ 5 =? Start by adding 4 to both sides 5+4=?

OpenStudy (anonymous):

9

pooja195 (pooja195):

good 3a≤9 divide both sides by 3 3/3=a a≤9/3 9/3=?

OpenStudy (anonymous):

3

pooja195 (pooja195):

good #1= a≤3

pooja195 (pooja195):

attempt #2 please

OpenStudy (anonymous):

what do you mean

pooja195 (pooja195):

Try: n/-3 + 5 > 4 By yourself

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

n/-3 + 5 > 4 = n>3

OpenStudy (anonymous):

is that it

pooja195 (pooja195):

Almost when you divide by a negative you need to flip the sign so n<3

OpenStudy (anonymous):

oh ok thankyou

pooja195 (pooja195):

yw :)